During March 2019, Annapolis Corporation recorded $42,800 of costs related to factory overhead. Alpha's overhead application rate is based on direct labor hours. The preset formula for overhead application estimated that $41,700 would be incurred, and 6,000 direct labor hours would be worked. During March, 4,250 hours were actually worked. Use this information to determine the amount of factory overhead that was (over) or under applied. (Round answers to the nearest whole dollar. Enter as a positive number if under applied. Enter as a negative number if over applied
Answer #1

Predetermined Overhead rate = Estimated overhead / Estimated direct labor hours

= 41,700/6,000 = 6.95 per labor hour

Applied overhead = 6.95 * 4250 = 29,538 (rounded)

Actual overhead = 42,800

So the overhead is

Underapplied by 13,262

Note : I have rounded some numbers. Comment if you face any issues